dan washes 3 cars in 2hrs. sue washes 3 cars in 1.5 hrs. how long to wash 3 cars together?
Answer by htmentor(1343)

dan washes 3 cars in 2hrs. sue washes 3 cars in 1.5 hrs. how long to wash 3 cars together?
==========================
Together they can wash 3 cars/2 hrs + 3 cars/1.5 hrs = 3/2 + 2 = 7/2 cars/hr
So it takes them 1/(7/2) = 2/7 hrs/car
So the time to wash 3 cars = (2/7)*3 = 6/7 hrs